OpenStudy (swati776):

The first term of a geometric progression is 25 and the sum of the first three terms is 61. Find the possible values of the common ratio. Given also that this progression has a sum to infinity, evaluate this sum.

OpenStudy (welshfella):

sum of n terms of a GS is a1. r^n - 1 ------ r - 1 where a1 = first term and r = common ratio

OpenStudy (faiqraees):

Sum of three numbers = first term(1-ratio^term number) ---------------------------- 1-ratio

OpenStudy (swati776):

Thank you, but i@m stuck here : 25r^3-61r+36=0

OpenStudy (welshfella):

so 61 = 25 ( r^3 - 1) ---------- r - 1

OpenStudy (welshfella):

yes a cubic equation like that is tricky to solve

OpenStudy (faiqraees):

Well use trial and error method to figure out one root, then divide this equation by (r-root) and solve the quadratic equation

OpenStudy (welshfella):

or use a graphical calculator if you have one!

OpenStudy (welshfella):

one solution is r = 1 but that cant be a valid value for r There are 2 more roots because its a cubic

OpenStudy (welshfella):

you could dive the equation by x -1 This would give you a quadratic

OpenStudy (welshfella):

- yes that is one way to go

OpenStudy (welshfella):

* by r - 1

OpenStudy (faiqraees):

@welshfella Thats what I said earlier. :)

OpenStudy (welshfella):

sorry - so you did!

OpenStudy (swati776):

thank you, i've found the answer however which value of r should i take into consideration, all 3 of them or 2 or 1?

OpenStudy (welshfella):

did you get 0.8 and -1.8?

OpenStudy (swati776):

yup

OpenStudy (welshfella):

ignore r = 1 as it would not be a GS in that case.

OpenStudy (welshfella):

Sum to infinity = a1 / (1 - r) so for r = 0.8 its 25 / ( 1 - 0.8)

OpenStudy (welshfella):

that formula is only valid for -1=< x <= 1

OpenStudy (swati776):

Ahhhhh i see, thanks

OpenStudy (welshfella):

What do you think about r = -1.8, Is that a possiblt value for the common ratio?

OpenStudy (swati776):

yes....

OpenStudy (welshfella):

It is but it doesn''t converge to a particular value so theres no sum to infinity.

OpenStudy (swati776):

ok, thanks again

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

its an oscillating series when r = -1.8 first few terms are 25, -45 , 81, -145.8